An investor plans to divide 200,000 between two investments. The first yields a certain profit of 10% while the second yields a profit with expected value of 18% and a standard deviation of 6%. If the investor divides the money equally between these two investments, find the mean and standard deviation of the total profit
1
Expert's answer
2020-12-23T05:06:40-0500

Linear combinations of Random Variables:

Let us consider a two random variable X and Y. let us define the linear combination of X and Y as follows.

"W = aX-bY"

Here, a and b are constants.

The mean or expected value of W is

"μ_W = a \times μ_X -b \times μ_Y"

The variance for W is as follows.

"σ_W^2 = a^2 \times σ_X^2 + b^2 \times σ_Y^2 -2 \times a \times b \times ρ(X,Y) \times σ_X \times σ_Y"

We are working on an investor plans. Let us define two random variables X and Y. X represent the first investment and Y represents second investments

"X = N(μ_X=0.1, σ_X^2 = 0) \\

Y = N(μ_Y=0.18, σ_Y^2 = 0.06)"

We are interested in computing the mean and variance of the total profit.

"P = 100000X + 100000Y"

We compute the mean and standard deviation of P by using previous formula

"a = 100000 \\

b = 100000 \\

μ_P = a \times μ_X + b \times μ_Y \\

= 100000 \times 0.1 + 100000 \times 0.18 \\

= 28000"

The mean of total profit is 28000.

"σ_W^2 = a^2 \times σ_X^2 + b^2 \times σ_Y^2 -2 \times a \times b \times ρ(X,Y) \times σ_X \times σ_Y \\



= (100000)^2 \times 0 + (100000)^2 \times (0.06)^2 -2 \times 100000 \times 100000 \times (0) \times 0.06 \\



= 36000000"

We know that the standard deviation is the positive square root of the variance, thus we have

"σ = \sqrt{36000000} \\

= 6000"

The standard deviation of total profit is 6000.